OpenStudy (anonymous):

let z1= 2 − 3i and z2 = −1 + 2i. compute: (z1+z2)^3 and (z1-conguate of z2)^-2

OpenStudy (experimentx):

z1+z2 = 2 − 3i −1 + 2i = 2 −1− 3i + 2i = 1 - i now we change 1 - i into a form where DeMoivre's theorem can be applied \[1 - i =\sqrt{2}(1/\sqrt{2} - i/\sqrt{2})\] = \[\sqrt{2}(\cos(-\pi/4) + i \sin (-\pi/4))\] (1-i)^3 = \[2^{3/2}(\cos(- 3 \pi/4) + \sin(-3\pi/4))\]

OpenStudy (zarkon):

simplifies nicely to -2-2i

OpenStudy (experimentx):

(z1-conguate of z2)^-2 = (2 − 3i −1 - 2i)^-2 = 1/(1-5i)^2 (1-5i)^2 simplify this .. ==> (1-25)+(-10)i => -24 - 10 i =-1/(24+10i) multiply both numerator and denominator by conjugate (24 - 10i) ... that should simplify the answer.